OpenStudy (anonymous):

For f(x)=2x+1 and g(x)=x^2-7, find (f+g)(x).

OpenStudy (anonymous):

A. x^2+2x-6 B. 2x^3-6 C. x^2+2x+8 D.2x^2-15

OpenStudy (shamil98):

x^2 -7 + 1 + 2x , you just combine the like terms. x^2 + 2x - 6

OpenStudy (anonymous):

ohh that totally makes sense! thank you @shamil98 :)

OpenStudy (shamil98):

You're welcome . :)

OpenStudy (anonymous):

@shamil98 for f(x)=3x+1 and g(x)=x^2-6, find (f-g)(x) how would i find that? the same way like you did above?

OpenStudy (shamil98):

Yes, but this time , you have to pay attention to (f -g ) okay? subtraction is different when the numbers are moved. So it would be : 3x + 1 - x^2 - 6 = -x^2 + 3x - 5

OpenStudy (shamil98):

Remember the properties of math. :)

OpenStudy (anonymous):

so basically instead of adding all of the numbers together, you combine them but subtract? @shamil98

OpenStudy (shamil98):

It says (F - G) when you are adding you can add either way (f + g ) or (g+f) the result would be different if it was (G - F )

OpenStudy (anonymous):

this is so confusing! :( lol ohhhh nevermind so i get the addition and subtracting one but what about multiplying them? like if it were f(x)=3x+1 and g(x)=x^2-6, find (f*g)(x)? @shamil98

OpenStudy (anonymous):

would i do (3x+1)(x^2-6)?

OpenStudy (shamil98):

It's multiplied so its distributed with the values, if you remember algebra 1 it comes handy here. so basically its like this : you multiply everything in one box with the other like 3x(x^2) and 3x(-6) 1(x^2) and 1(-6) then you combine the like terms.

OpenStudy (shamil98):

3x^3 - 18x + x^2 - 6 3x^3 +x^2 - 18x -6 is the answer.

OpenStudy (anonymous):

oh okay, so would the answer to that be 3x^3+x^2-18x-6?

OpenStudy (anonymous):

oh hehe just saw your answer, woohoo! I got it. thank you soo soo much i really appreciate it!! :)

OpenStudy (shamil98):

Yeah, no problem.